OpenStudy (lovelyharmonics):

[03.09]Calculate the percent composition of calcium in calcium hydroxide. 16% 33% 54% 70%

OpenStudy (lovelyharmonics):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Can you make the molecular formula for calcium hydroxide? :3 I told you hydroxide is OH and has a -1 charge. Calcium has a +2 charge in compounds.

OpenStudy (lovelyharmonics):

How does this have to do with the formula? It's percentages

OpenStudy (ipwnbunnies):

Yeah, percent composition. How many calcium atoms make up a molecule of calcium hydroxide. The molecular formula will tell us. :3

OpenStudy (lovelyharmonics):

I vote that its 54% .-. It sounds like it should be half

OpenStudy (ipwnbunnies):

Why can't you just make the molecular formula? .-.

OpenStudy (ipwnbunnies):

Calcium has a 2+ charge. Hydroxide, a 1- charge. The molecule has to be 0 charge. So, we need another ion of hydroxide to make this molecule. Ca(OH)2 is the molecular formula for calcium hydroxide. Oh crap, idk if this is atoms or mass. .-.

OpenStudy (ipwnbunnies):

Ok, yah, it is related to mass. Can you find the mass of a molecule of Ca(OH)2? :33

OpenStudy (lovelyharmonics):

Ca-40.08 O-16H-1.008

OpenStudy (ipwnbunnies):

Mhm. And there are 2 atoms of oxygen and hydrogen per molecule.

OpenStudy (lovelyharmonics):

So 32 and 2

OpenStudy (ipwnbunnies):

Yas. So the mass of one molecule is 40.08 + 32 + 2 amu = 74 amu Since there's one atom of calcium, the mass of calcium is just 40 amu. Percent composition = 40/74 * 100% = 54%

OpenStudy (lovelyharmonics):

Yay I was right c:

OpenStudy (somy):

Ca(OH)2 get full molar mass molar mass of the whole thing - 100% molar mass of Ca- X X= (molar mass of Ca* 100) / molar mass if the whole thing meaning Ca(OH)2

OpenStudy (ipwnbunnies):

Somy. ;3

OpenStudy (somy):

can we do it mathematically? Bunnies <3 q is for u

OpenStudy (somy):

i mean i didn't think of formula or anything

OpenStudy (ipwnbunnies):

I did do it mathematically, I thought. :3

OpenStudy (somy):

lmao i didn't think what i said lol